What is componendo and dividendo, how do we apply this in different fields of mathematics?

Grade:11

2 Answers

Himanshu
7 years ago
Componendo and dividendo are just a part of the mathematics calculation.
Componendo and dividendo are greek words. According to it,
If a/b = c/d , then (a+b) / (a-b) = (c+d) / (c-d).
Let me give you proof for the above statement.
Add 1 to both sides of the equation i.e., a/b + 1 = c/d + 1
(a+b) / b = (c+d) / d                      ….i)
Now, Subtract 1 from both sides of the equation i.e., a/b - 1 = c/d - 1
(a-b) / b = (c-d) / d                        ….ii)
Dividing eq. i) by ii). We get,
(a+b) / (a-b) = (c+d) / (c-d)
We use componendo and dividendo just to speed up the calculation part of the mathematics.
 
Kartik pandey
6 years ago
If a/b=c/d then we can say that a:b=c:d=k(universally recognised as costant).If a/b=c/d then a+b/a-b=c+d/c-d.By tranfering b on the other side we can say that a=kb,c=kd.A.T.Q kb+b/kb-b=kd+d/kd-dLets take LHS first:Lets take b common now we can say that b(k+1)/b(k-1) cancel out b fron numerator and denominator we obtain (k+1)/(k-1)RHS:kd+d/kd-d take d as common we can say that d(k+1)/d(k-1) .Cancel out d fron both numerator and denominator we obtain (k+1)/(k-1).LHS = (k+1)/(k-1)= RHS =(k+1)/(k-1).We conclude that LHS=RHS QED.